OpenStudy (anonymous):

Dora drew JKL and MNP so that K N, JK = 6, MN = 18, KL = 4, and NP = 12. Are JKL and MNP similar? If so, identify the similarity postulate or theorem that applies.

OpenStudy (anonymous):

\[\frac{ JK }{ KL }=\frac{ MN }{NP } \frac{ 6 }{ 4 }=\frac{ 18 }{ 12}\] Now u cross multiply. if both numbers are equal, they are similar. In this case, they are similar. 6 x 12 = 4 x 18 72=72

OpenStudy (anonymous):

though the two corresponding sides of the triangles mentioned are proportional (proportionality being 1:3 ) but they need not be similar for similarity either the third corresponding sides must be proportional or the included angle between the two proportional sides must be equal

OpenStudy (anonymous):

So is it SAS (Side angle side)

OpenStudy (anonymous):

Or SSS (Side side side)

OpenStudy (anonymous):

Um, i was never good with that. I dont want to mislead you. lol

OpenStudy (anonymous):

Ok thnks anyways it helped

OpenStudy (anonymous):

but... if i had to choose i would say SSS because we know 2 sides are similar/proportional. So we can conclude the same for the 3rd side.... and the problem doesn't say anything regarding an angle so i would not choose SSA or a s s backwards, lol.

OpenStudy (anonymous):

Thnk you very much :D

OpenStudy (anonymous):

No problem

OpenStudy (jhannybean):



OpenStudy (jhannybean):

\[\large \frac{JK}{MN}=\frac{KL}{NP} =\frac{6}{18}=\frac{4}{12} \]Creates a \(1:3\) ratio...hm.. Do you have options?

OpenStudy (anonymous):

I got it thank you anyways @FilthyMcNasty was correct It was similar-SSS

OpenStudy (jhannybean):

Oh I see. if two sides in a triangle are congruent, then the 3rd side must be congruent as well :) Cool.

OpenStudy (anonymous):

Yes thank you tho

OpenStudy (jhannybean):

NP ^^

OpenStudy (anonymous):

I have one more question how do I simplify 3x=x+10

OpenStudy (jhannybean):

bring the x's to one side and your constants on the other. then solve. Start by subtracting x to both sides. 3x - x = -x +x + 10 Simplify the equation 2x = 10. Divide both sides by 2. 2x =10/2 x= ?

OpenStudy (anonymous):

Ok thnk you